Simplify combining like terms: 
21b – 32 + 7b – 20b
Simplify combining like terms: 
21b – 32 + 7b – 20b

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

When term have the same algebraic factors, they are like terms.
Then,
= (21b + 7b–20b)–32
= b (21 + 7–20)–32
= b (28 - 20)–32
= b (8) - 32
= 8b - 32

Thanks